# Robust Point Matching with Distance Profiles

### v5

@@ Abstract
We show the outlier robustness and noise stability of practical matching procedures based on distance profiles. Although the idea of matching points based on invariants like distance profiles has a long history in the literature, there has been little understanding of the theoretical properties of such procedures, especially in the presence of outliers and noise. We provide a theoretical analysis showing that under certain probabilistic settings, the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings. Lastly, we extend the concept of distance profiles to the abstract setting and connect the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound, with a new sample complexity result derived based on the properties of distance profiles. This paper contributes to the literature by providing theoretical underpinnings of the matching procedures based on invariants like distance profiles, which have been widely used in practice but have rarely been analyzed theoretically.
@@ 1 Introduction
Matching is one of the most fundamental tasks of modern data science, which can be found in numerous applications. For instance, the goal of shape correspondence problems [37] in computer graphics is to match corresponding parts of different geometric shapes, often obtained by 3D scanners or imaging equipment used for Magnetic Resonance Imaging (MRI) or Computed Tomography (CT). Graph matching aims to find corresponding nodes between two or more graphs representing structured data, such as neuronal connectivity in human brains [35] or social networks [6].
@@
A formal mathematical framework for matching [26] conceptualizes objects as metric measure spaces, which are essentially probability measures defined on suitable metric spaces. In practice, of course, objects are not fully available and thus approximated by suitable point clouds, say, $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ on $\mathbb{R}^{d}$, respectively, and the practical goal is to match these point clouds. Quadratic matching is one of the most standard approaches to this problem, which formulates and solves a quadratic program based on pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$. When $n=m$, quadratic matching finds a one-to-one correspondence that aligns the pairwise distances as closely as possible, which is formulated as a Quadratic Assignment Problem (QAP) [21, 24]. When $n\neq m$, the notion of correspondence is relaxed to soft assignments or couplings, leading to the formulation of the Gromov-Wasserstein distance [26], which has been increasingly used in applications, such as word alignment [2], statistical inference [7, 41], and generative modeling [9, 19]. On the practical side, however, quadratic matching requires solving a quadratic program that is NP-hard in general, and thus needs to rely on several methods, such as convex relaxation [1], semidefinite relaxation [42, 20], spectral relaxation [23], or entropic regularization [32, 34, 31].
@@
Matching procedures based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n},\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are invariant to rigid transformations, which is desirable in practice. For instance, in imaging problems, the images of the objects of interest are often obtained from different angles and thus need to be matched up to rigid transformations. However, such procedures are difficult to analyze theoretically and the literature has been mostly studying theoretical guarantees for other matching procedures using $\|X_{i}-Y_{j}\|_{2}$’s, which aim to match the points based on the proximity. [11, 22, 40] establish theoretical results on recovering the ground truth matching for the procedures based on $\|X_{i}-Y_{j}\|_{2}$’s under some noise, and [18, 28, 17] show matching guarantees in the presence of outliers. Although Theorem 2 of [40] concerns a procedure using the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, it requires solving a nontrivial minimization over permutations, which is computationally challenging, and the result does not consider the presence of outliers.
@@
Motivated by the above challenges, this paper theoretically establishes outlier robustness and noise stability of a matching procedure based on the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, which is easily implementable in practice. The proposed procedure associates each point with the set of distances from it to other points, namely, $X_{i}$ and $Y_{j}$ are associated with the sets of distances to the other points $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$, which we call the distance profiles of $X_{i}$ and $Y_{j}$, respectively. The matching procedure is to find for each $X_{i}$, the point $Y_{j}$ having the most similar distance profile to that of $X_{i}$, where the similarity is measured by a suitable quantity between the sets $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$. The idea of associating a point with suitable invariants based on the distances to other points from it has a long history in the object matching literature, which includes shape contexts [4, 33], integral invariants [25], invariant histograms [8], and the local distribution of distances [26]. However, theoretical results on robustness of such methods are rarely found in the literature. In the context of random graph matching, [12, 13] propose and analyze a degree profile matching procedure based on the adjacency matrices $(A_{i\ell})_{i,\ell=1}^{n}$ and $(B_{j\ell})_{j,\ell=1}^{n}$ of certain correlated random graphs. However, their analysis focuses on random graph models, such as Erdös-Rényi and correlated Wigner models, where the edges are independently generated, namely, $(A_{i\ell})_{i,\ell=1}^{n}$ are mutually independent, which is substantially different from our setting where the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ are not mutually independent.
Our contribution
We show the outlier robustness and noise stability of practical matching procedures based on distance profiles. The idea of matching points based on invariants like distance profiles has a long history in the literature, proposed and rediscovered multiple times under different names and contexts [4, 33, 25, 8, 26, 12, 13]. However, there has been little understanding of the theoretical properties of such procedures, especially in the presence of outliers and noise. We provide a theoretical analysis showing that under certain probabilistic settings, the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We check the performance of the proposed method using a stylized data example and provide simulation studies to complement the theoretical findings. Lastly, we extend the concept of distance profiles to the abstract setting and connect the proposed matching procedure to the Gromov-Wasserstein distance and its lower bound [26], with a new sample complexity result derived based on the properties of distance profiles. As a result, we contribute to the literature by providing theoretical underpinnings of the matching procedures based on invariants like distance profiles, which have been widely used in practice but have rarely been analyzed theoretically.
@@
1.1 Distance Profile Matching
We first describe the distance profile matching procedure that we propose to study. As mentioned earlier, we build on the principle of matching points using invariants like distance profiles, which has a long history in the literature [4, 33, 25, 8, 26, 12, 13]. The idea is to associate each point with the set of distances from it to other points, which we call the distance profile, and match the points based on the similarity of the distance profiles. While there are many different procedures incorporating this idea, we are particularly inspired by a simple and easily implementable procedure called degree profile matching [12] that was proposed to match random graphs. We properly modify it to match point clouds, which we summarize in Algorithm 1. We call this procedure the distance profile matching.
First, notice that Algorithm 1 is implementable without knowing the points $X_{i}$’s and $Y_{j}$’s as long as the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are provided. The discrepancy measure $\mathcal{D}$ quantifies the similarity between the distance profiles. Therefore, for each $X_{i}$, the procedure (1.2) finds $Y_{j}$ whose distance profile $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$ is closest to $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$, the distance profile of $X_{i}$, under $\mathcal{D}$.
The output $\pi$ of Algorithm 1 represents the matching: $X_{i}$ is matched to $Y_{\pi(i)}$. Note that $\pi$ may not be injective or surjective, meaning that some points may not be matched or some points may be matched to multiple points. [^1] When $n=m$, Algorithm 1 can be modified to output a permutation (one-to-one correspondence); see Section 3. The motivation behind this is simple: (1) in the presence of outliers, some points can be ignored or matched arbitrarily, and (2) similar points can be matched to the same point. The underlying objects that the points $X_{i}$’s and $Y_{j}$’s approximate, say, probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{d}$, respectively, are not identical in practice. Usually, they coincide up to several outlier parts, which we can formally state as $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$ for some $\lambda\in(0,1)$ and $\rho,\mu_{o},\nu_{o}\in\mathscr{P}(\mathbb{R}^{d})$. Here, $\lambda\rho$ represents the common part of $\mu$ and $\nu$ that we want to match, while $(1-\lambda)\mu_{o}$ and $(1-\lambda)\nu_{o}$ are the outlier parts of $\mu$ and $\nu$, respectively. We are mainly interested in matching $X_{i}$’s and $Y_{j}$’s from the common part, while ignoring or arbitrarily matching the points from the outlier parts.
The threshold $\rho$ is an optional parameter that can be used to obtain the subset $I\subset[n]$ containing the indices $i$’s such that the discrepancy $D(i,\pi(i))$ between the distance profiles of $X_{i}$ and its match $Y_{\pi(i)}$ is below $\rho$. Providing a small threshold leads to a more conservative matching that only focuses on the points that are closely matched, which can be used for other downstream tasks. For instance, one can solve the orthogonal Procrustes problem using the matched pairs $\{(X_{i},Y_{\pi(i)})\}_{i\in I}$ to find the optimal rotation and translation aligning them as closely as possible, which theoretically works with just $|I|=d$ pairs in $\mathbb{R}^{d}$ but practically requires more pairs, say, $|I|=O(d)$, due to potential noise or degeneracy. The obtained transformation can then be used as an initialization of the registration procedures, such as the Iterative Closest Point (ICP) algorithm [5], to find the optimal transformation that aligns the whole point clouds.
In this paper, we mainly use the Wasserstein-1 distance $W_{1}$ as the discrepancy measure so that $D(i,j)$ in (1.2) becomes
$W(i,j):=W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}}\right)\quad\forall(i,j)\in[n]\times[m].$ (1.1)
The Wasserstein-1 distance between two sets of points in $\mathbb{R}$ admits a closed-form expression based on the corresponding cumulative distribution functions, which can be computed efficiently [30]. In this case, the computational complexity of Algorithm 1 is computed as follows. For Step 1, assuming $n\geq m$, computing the Wasserstein-1 distances between two sets of points in $\mathbb{R}$ for $nm$ pairs takes $O(n^{3}\log(n))$ in total. Meanwhile, Step 2 takes $O(nm\log(m))$ as it requires sorting $m$ numbers $n$ times. Therefore, the overall complexity of Algorithm 1 is $O(n^{3}\log(n))$ assuming $n\geq m$, which is nearly cubic in the number of points $n$; however, as the input dimension of the algorithm is $n^{2}$, we can say that the complexity scales nearly as the input dimension $n^{2}$ raised to the power of $\frac{3}{2}$.

### v6

@@ Abstract
Computational difficulty of quadratic matching and the Gromov-Wasserstein distance has led to various approximation and relaxation schemes. One of such methods, relying on the notion of distance profiles, has been widely used in practice, but its theoretical understanding is limited. By delving into the statistical complexity of the previously proposed method based on distance profiles, we show that it suffers from the curse of dimensionality unless we make certain assumptions on the underlying metric measure spaces. Building on this insight, we propose and analyze a modified matching procedure that can be used to robustly match points under a certain probabilistic setting. We demonstrate the performance of the proposed methods using simulations and real data applications to complement the theoretical findings. As a result, we contribute to the literature by providing theoretical underpinnings of the matching procedures based on distance invariants like distance profiles, which have been widely used in practice but rarely analyzed theoretically.
@@ 1 Introduction
Matching is one of the most fundamental tasks of modern data science, which can be found in numerous applications. For instance, the goal of shape correspondence problems [van2011survey, sahilliouglu2020recent] in computer graphics is to match corresponding parts of different geometric shapes, often obtained by 3D scanners or imaging equipment used for Magnetic Resonance Imaging (MRI) or Computed Tomography (CT). Graph matching aims to find corresponding nodes between two or more graphs representing structured data, such as neuronal connectivity in human brains [sporns2005human] or social networks [borgatti2009network].
@@
One formal mathematical framework for matching [memoli2011gromov] is to conceptualize objects as metric measure spaces, which are essentially probability measures defined on suitable metric spaces. In practice, of course, objects are not fully available and thus approximated by suitable point clouds, say, $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ on metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$, respectively, and the practical goal is to match these point clouds. Quadratic matching is one of the most standard approaches to this problem, which formulates and solves a quadratic program based on pairwise distances $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{i,\ell=1}^{n}$ and $\{d_{\mathcal{Y}}(Y_{j},Y_{\ell})\}_{j,\ell=1}^{m}$. When $n=m$, quadratic matching finds a one-to-one correspondence that aligns the pairwise distances as closely as possible, which is formulated as a Quadratic Assignment Problem (QAP) [koopmans1957assignment, loiola2007survey]. When $n\neq m$, the notion of correspondence is relaxed to soft assignments or couplings, leading to the formulation of the Gromov-Wasserstein distance [memoli2011gromov], which has been increasingly used in applications, such as word alignment [alvarez2018gromov], statistical inference [brecheteau2019statistical, weitkamp2022distribution], and generative modeling [bunne2019learning, hur2024reversible]. On the practical side, however, quadratic matching requires solving a quadratic program that is NP-hard in general, and thus is approximated by several methods, such as convex relaxation [aflalo2015convex], semidefinite relaxation [zhao1998semidefinite, kezurer2015tight], spectral relaxation [leordeanu2005spectral], or entropic regularization [rangarajan1999convergence, solomon2016entropic, peyre2016gromov].
@@
Matching procedures based on pairwise distances $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{i,\ell=1}^{n}$ and $\{d_{\mathcal{Y}}(Y_{j},Y_{\ell})\}_{j,\ell=1}^{m}$ are invariant to isometries, which is desirable in practice. For instance, in imaging problems, the images of the objects of interest are often obtained from different angles and thus need to be matched up to rigid transformations. However, such procedures are difficult to analyze theoretically and the literature has been mostly studying theoretical guarantees for other matching procedures based on $\|X_{i}-Y_{j}\|_{2}$’s when $(\mathcal{X},d_{\mathcal{X}})=(\mathcal{Y},d_{\mathcal{Y}})$ is the standard Euclidean space. In this setting, the point clouds lie in the same space and are often viewed as perturbations of one another, where one aims to match the points based on the proximity. [collier2016minimax, kunisky2022strong, wang2022random] establish theoretical results on recovering the ground truth matching for the procedures based on $\|X_{i}-Y_{j}\|_{2}$’s under some noise, and [galstyan2022optimal, minasyan2023matching, galstyan2023optimality] show matching guarantees in the presence of outliers. Although Theorem 2 of [wang2022random] concerns a procedure using the pairwise distances in the Euclidean setting, it requires solving a nontrivial minimization over permutations, which is computationally challenging, and the result does not consider the presence of outliers.
@@
Motivated by the above challenges, this paper theoretically analyzes matching procedures based on the distances $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{i,\ell=1}^{n}$ and $\{d_{\mathcal{Y}}(Y_{j},Y_{\ell})\}_{j,\ell=1}^{m}$, which are easily implementable in practice. The idea is to associate each point with the set of distances from it to other points, namely, $X_{i}$ and $Y_{j}$ are associated with the sets of distances to the other points $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{\ell=1}^{n}$ and $\{d_{\mathcal{Y}}(Y_{j},Y_{\ell})\}_{\ell=1}^{m}$, which are the distance profiles of $X_{i}$ and $Y_{j}$, respectively. The matching procedure is to find for each $X_{i}$, the point $Y_{j}$ having the most similar distance profile to that of $X_{i}$, where the similarity is measured by a suitable quantity between the sets $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{\ell=1}^{n}$ and $\{d_{\mathcal{Y}}(Y_{j},Y_{\ell})\}_{\ell=1}^{m}$. The idea of associating a point with suitable invariants based on the distances to other points from it has a long history in the object matching literature, which includes shape contexts [belongie2002shape, ruggeri2008isometry], integral invariants [manay2006integral], invariant histograms [brinkman2012invariant], the local distribution of distances [memoli2011gromov], and other related concepts [gortler2019generic, norelli2023asif]. However, theoretical results on robustness of such methods are rarely found in the literature. In the context of random graph matching, [ding2021efficient, ding2023efficiently] propose and analyze a degree profile matching procedure based on the adjacency matrices $(A_{i\ell})_{i,\ell=1}^{n}$ and $(B_{j\ell})_{j,\ell=1}^{n}$ of certain correlated random graphs. However, their analysis focuses on random graph models, such as Erdös-Rényi and correlated Wigner models, where the edges are independently generated, namely, $(A_{i\ell})_{i,\ell=1}^{n}$ are mutually independent, which is substantially different from our setting where the distances $\{d_{\mathcal{X}}(X_{i},X_{\ell})\}_{i,\ell=1}^{n}$ are not mutually independent.
Organization
Section 2 presents mathematical background on the Gromov-Wasserstein distance and its computational alternative based on the distance profile, together with the new sample complexity result in Section 2.3. Motivated by the insights built in Section 2, Section 3 proposes and studies the partial matching algorithm based on distance profiles, where we formally show its robustness to noise and outlier components, together with the noise stability analysis for the one-to-one matching procedure. Section 4 demonstrates the performance of the proposed methods using simulations. All technical proofs, lemmas, extensions are deferred to Appendices A, B, C, respectively, to streamline the main text.
@@
2 Foundations: Distance Profile Distance and Sample Complexity
As mentioned in Section 1, computational difficulty of the Gromov-Wasserstein (GW) distance serves as a major obstacle in practical applications, which has led to the development of various approximation and relaxation schemes. One of such methods, proposed as a tractable lower bound to the GW distance by [memoli2011gromov], utilizes the notion of distance profiles. We delineate the mathematical foundations of the distance profile in connection with the GW distance and see how the distance based on it, which we call the distance profile distance (DPD), avoids the computational difficulty of the GW distance. Then, we derive new results on the sample complexity of the DPD and discuss the limitations and potentials of the DPD as a matching method under certain probabilistic settings, which motivates the distance profile matching procedure in the next section.
2.1 Background: Gromov-Wasserstein Distance and Computational Difficulty
Following [memoli2011gromov], we build on the mathematical framework of object matching based on metric measure spaces. The Gromov-Wasserstein (GW) distance is a generalization of quadratic matching between point clouds to abstract metric measure spaces. The GW distance is defined based on the notion of couplings, which are probability measures on the product space of two metric measure spaces that preserve the marginal distributions of the two spaces. Couplings represent the correspondences between the points in the two metric measure spaces, and the GW distance finds the optimal coupling that best aligns the pairwise distances in the two spaces. Formal definitions are as follows.
Definition 1.
Definition 2.
Definition 3.
The objective function on the right-hand side of (2.1) is quadratic in the variable $\gamma$ and is not convex in general, which makes the minimization computationally challenging. The following discrete case illustrates the computational difficulty of the GW distance in the case of point clouds, which is a common setting in practice.
Example 1 (Discrete Case).
[…]
where $\Pi_{n,m}=\{\gamma\in\mathbb{R}_{+}^{n\times m}:\sum_{j=1}^{m}\gamma_{ij}=\frac{1}{n}~\forall i\in[n]~\text{and}~\sum_{i=1}^{n}\gamma_{ij}=\frac{1}{m}~\forall j\in[m]\}$ represents the set of all couplings between $\hat{\mu}$ and $\hat{\nu}$ as nonnegative matrices—called coupling matrices—of which the sum of each row (resp. column) is $\frac{1}{n}$ (resp. $\frac{1}{m}$). We again observe that the objective function of (2.2) is quadratic in the variable $\gamma$, which is NP-hard to solve in general.
2.2 Distance Profile Distance
Due to the aforementioned computational difficulty of the GW distance, various approximation and relaxation schemes have been proposed in the literature. [memoli2011gromov] proposes several relaxations to the GW distance, which provide lower bounds to the GW distance. One of the three lower bounds proposed in [memoli2011gromov], which he named the third lower bound based on the ordering, is based on the idea of local distribution of distances, which we generalize as distance profiles below.
Definition 4.
In a nutshell, for each $x\in\mathcal{X}$, the distance profile $\mu_{x}$ is the distribution of the distances from $x$ to the points in $\mathcal{X}$ with respect to the measure $\mu$. This leads to a natural matching procedure as follows: given two metric measure spaces $(\mathcal{X},d_{\mathcal{X}},\mu)$ and $(\mathcal{Y},d_{\mathcal{Y}},\nu)$, we find a coupling that best aligns the distance profiles, namely, associate $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ if the distance profiles $\mu_{x},\nu_{y}\in\mathscr{P}(\mathbb{R})$ are similar. Here, similarity is measured by any suitable discrepancy on $\mathscr{P}(\mathbb{R})$; we use the Wasserstein distance as a natural choice, which leads to the following definition.
Definition 5.
The distance profile distance (DPD) of order $p\geq 1$ between metric measure spaces $(\mathcal{X},d_{\mathcal{X}},\mu)$ and $(\mathcal{Y},d_{\mathcal{Y}},\nu)$ is defined as follows:
$\mathfrak{T}_{p}(\mu,\nu):=\left(\inf_{\gamma\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}W_{p}^{p}(\mu_{x},\nu_{y})\,\mathrm{d}\gamma(x,y)\right)^{1/p}.$ (2.3)
Remark 1.
In Definition 5, while we use $W_{p}^{p}$ to measure the discrepancy between the distance profiles $\mu_{x},\nu_{y}\in\mathscr{P}(\mathbb{R})$, any suitable discrepancy measure on $\mathscr{P}(\mathbb{R})$ can be used in place of $W_{p}^{p}$. The choice $W_{p}^{p}$ leads to a nice metric property, namely, $\mathfrak{T}_{p}$ is a pseudometric on the set of metric measure spaces, and also naturally connects to the Gromov-Wasserstein distance. Indeed, one can show that $\mathfrak{T}_{p}\leq\mathrm{GW}_{p}$. See [memoli2011gromov] and [memoli2022distance] for technical details.
Unlike the GW distance, the objective function on the right-hand side of (2.3) is linear in the variable $\gamma$, which makes the minimization computationally tractable. The following discrete case illustrates the computational advantage of the distance profile distance in the case of point clouds.
Example 2.
[…]
where $\Pi_{n,m}$ is as defined in Example 1. Computational advantage of $\mathfrak{T}_{p}$ over $\mathrm{GW}_{p}$ is clear: $\mathfrak{T}_{p}$ can be computed by solving a linear program, while $\mathrm{GW}_{p}$ requires solving a non-convex quadratic program. Unlike (2.2), the minimization (2.4) is a linear program that is solvable by polynomial-time algorithms.
2.3 Sample Complexity: Challenges and Opportunities
[…]
Setting
[…]
2.3.1 General Case: Curse of Dimensionality
We first observe that
where the last term can be shown to converge at the rate of $O(n^{-1/2})$ as stated below.
Theorem 1.
where $C$ is an absolute constant that is independent of $n$, $d$, and $\mu$.
Remark 2.
For each fixed $x\in\mathcal{X}$, as $\mu_{x}$ is a probability measure on $\mathbb{R}$, one can deduce from the standard results [dudley1969speed, fournier2015rate] that $W_{1}(\mu_{x},\hat{\mu}_{x})$ enjoys a fast rate of convergence in terms of the sample size $n$, which is $O(n^{-1/2})$. Theorem 1 shows that taking the supremum over $x\in\mathcal{X}$ does not slow down the rate of convergence. We present the extension of Theorem 1 to general metric measure spaces in Appendix C.1.
[…]
Proposition 1.
[…]
2.3.2 Faster Rate by Invariance and Effective Dimension
It turns out that when $\mu$ has a certain structure, we can derive a faster rate of convergence for (*). Namely, if $\mu$ is a mixture of rotationally invariant distributions like a Gaussian mixture, then the effective dimension of $\mu$ is the number of the components, which can be much smaller than the ambient dimension $d$. The key is that $W_{1}(\mu_{x},\mu_{x^{\prime}})$, the distance between distance profiles, captures this structure and thus is determined by the effective dimension, rather than the ambient dimension.
Proposition 2.
[…]
Theorem 2.
where $C$ is an absolute constant that is independent of $n$, $d$, and $\mu$.
Summary and Insights
The above results provide insights into how to use distance profiles for matching. Due to the curse of dimensionality, it is impractical to estimate the distance profile distance $\mathfrak{T}_{1}$ between high-dimensional metric measure spaces by random samples. Hence, it is more suitable to use and study $\mathfrak{T}_{1}$ under a different setting than the sampling scenario; the next section conducts such a study based on the model with fixed high-dimensional locations corrupted by noise. Meanwhile, we have seen that a better rate can be achieved under a mixture model with rotationally invariant components. The high-level idea is that when such a structure is present, points from the same component are treated as the same point so that the effective dimension is essentially the number of components. This is pertinent to partial matching, where we want to allow similar points from one object to be matched to the same point in another object, not necessarily requiring a one-to-one correspondence or a coupling. In the next section, we study the distance profile matching procedure that allows for such flexibility, with a robustness property.
3 Distance Profile Matching and Its Theoretical Guarantees
This section introduces and analyzes the distance profile matching algorithms, inspired by the insights on the distance profile distance $\mathfrak{T}_{1}$ in the previous section. We have learned that using $\mathfrak{T}_{1}$ directly for matching in the random sampling setting can be impractical. Hence, we propose a flexible matching procedure—Algorithm 1—modified from $\mathfrak{T}_{1}$ that allows for partial matching and derive its robustness property. Meanwhile, Algorithm 2 is a special case of $\mathfrak{T}_{1}$ to find a one-to-one correspondence, which we analyze under a fixed-location model with noise, deviating from the random sampling setting where $\mathfrak{T}_{1}$ is impractical.
3.1 Distance Profile Matching Algorithms
3.1.1 Distance Profile Matching Algorithm for Partial Matching